OpenStudy (anonymous):

how do i use macluerins expansion on (e^2x -1)/(e^2x +1)

OpenStudy (anonymous):

just to clarify \[e ^{2x}-1 \over e ^{2x}+1 \]

OpenStudy (anonymous):

i need the terms to include x to power 5

OpenStudy (anonymous):

now i know \[e^{2x}=1+ 2x + {(2x)^2 \over2!} + ...\]

OpenStudy (anonymous):

but i'm not sure how that helps with this

OpenStudy (anonymous):

you are on the right track

OpenStudy (anonymous):

problem is i'll end up with the terms of the expansion in numerator and denominator

OpenStudy (anonymous):

plus i've tried to simplify to \[1- {2\over e^{2x} + 1}\]

OpenStudy (anonymous):

i've also tried directly differentiating the whole expression but it rapidly gets complex after 2 levels of differentiation

OpenStudy (anonymous):

\[\frac{1+ 2x + {(2x)^2 \over2!} + ... -1}{1+ 2x + {(2x)^2 \over2!} + ... +1}\] \[\frac{ 2x + {(2x)^2 \over2!} + ... }{2+ 2x + {(2x)^2 \over2!} + ... }\]

OpenStudy (anonymous):

apparently the answer is \[\approx x - {x^3\over 3} +{ 2x^5 \over15}\]

OpenStudy (anonymous):

i got as far as you imran but nowhere near that answer